[Hallc_running] [New Logentry] RC Daily Update

pooser at jlab.org pooser at jlab.org
Mon Nov 19 10:15:01 EST 2018

Logentry Text:
RC report:
We began yesterday shift collecting data for kinematic group 14.  At 09:30 we initiated a BCM calibration which was successfully completed at ~11:00 (https://logbooks.jlab.org/entry/3628176) at which point we continued on with group 14.  At ~05:45 we completed kinematic group 14 and switched polarity on the SHMS (https://logbooks.jlab.org/entry/3628389) in order to return to the lowest z-setting in group 13.

Throughout the weekend beam was quite reliable (https://logbooks.jlab.org/entry/3628473) and ~3 kinematic groups were completed.  Currently we have completed kinematic group 13-10 (LD2) and have begun 13-11 (dummy).  It is estimated that we will complete 13-12 (LH2) around the start os swing.  One group 13 is complete we will configure the spectrometers for group 16.

Notices in effect for shift crews:
Shift leaders, please fill out the check list once per shift!
It is no longer required to fill out the standard.kinematics file online, it is now done automagically
Split the large desired charge between multiple runs (~1 hour per run)
Listed currents are estimates. For runs at less than 60uA, increase current until SHMS singles rates are between 400k-500k.
How-to power cycle PLC I/O chassis: https://logbooks.jlab.org/entry/3625862, https://logbooks.jlab.org/entry/3625900

Planned accelerator configuration changes:
Tuesday, Nov 20: CHL UPS battery change, should be noninvasive however, induce CHL anomaly is possible
Wednesday, Nov 21: Hall A goes offline and MCC wants to conduct high current beam study of energy spread at 40 uA.  It is unclear that we will be effected
Monday, Nov 26: Energy change to 1.86 GeV/pass, we will be moved over to 500 MHz laser.

Planned accesses to the hall:
Monday, Nov 26: aerogel change and HGC work needs to be coordinated with Walter K., we must be in restricted access for this work

Special requests:
Harp scan at 3H07A and 3H07B with raster off to address possible issues with harp (https://logbooks.jlab.org/entry/3624463)
Vijay Kumar (Regina group) has requested a walkthrough during the next access
Jay B. would like to conduct another beam steering test (5 minute procedure) during the next polarity change.  He requests to only be contacted if the polarity change occurs during the hours of 0630 and 1930.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3628475
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20181119/9ecbc489/attachment.html>

More information about the Hallc_running mailing list